# ConsRec: Denoising Sequential Recommendation through User-Consistent Preference Modeling

### v1

@@ 5.2 Ablation Study
Figure 3: Effectiveness of ConsRec in Learning Item Representations for Denoising. Figure 3(a) and Figure 3(b) illustrate the item similarities among $M_{\text{Filter}}$-retained items and the target items.
(a) Raw User-Item Interaction Trained $M_{\text{Rec}}$.
(b) $M_{\text{Filter}}$-Retained Items Trained $M_{\text{Rec}}$.
Figure 4: Performance of Recommendation Models Using Different Item Denoising Methods.
5.3 Effectiveness of Different Item Representation Strategies in Denoising
To assess the effectiveness of the denoising module in ConsRec, we conduct a comparative analysis involving various item representation strategies. Specifically, we compare the full ConsRec model with its variant without the denoising module (ConsRec w/o Filter), as well as versions of ConsRec equipped with different item representation methods within the denoising module. We implement the denoising module $M_{\text{Filter}}$ using three approaches: an ID-based recommendation model, Bert4Rec Sun et al. (2019), and a dense retriever, BGE-Base Xiao et al. (2023), which serve as baselines for evaluating representation effectiveness.
@@ 5.3 Effectiveness of Different Item Representation Strategies in Denoising
As illustrated in Figure 3, we analyze the text similarity among the items retained by $M_{\text{Filter}}$ using the BGE-Base model to compute pairwise similarity scores. In Figure 3(a), we report the average similarity among the retained items using different denoising methods. The results show that all methods improve the internal similarity of the retained items. Notably, when using ConsRec, the retained user-interacted items achieve the highest average similarity, indicating stronger coherence among these retained items. Furthermore, Figure 3(b) presents the average similarity between the $M_{\text{Filter}}$-retained user-interacted items and the ground-truth target item. Among all methods, ConsRec yields the highest similarity, demonstrating its effectiveness in retaining items that are semantically aligned with the target item that users needed.
@@
As illustrated in Figure 4, we evaluate the effectiveness of various denoising methods across different recommendation systems. Figures 4(a) and 4(b) report the performance of the recommendation module $M_{\text{Rec}}$ when trained on either the raw user-item interaction set $\mathcal{H}$ or the denoised subset $\mathcal{H}_{\text{Filter}}$ retained by $M_{\text{Filter}}$. In both cases, ConsRec consistently outperforms other item representation models, with particularly strong gains at larger ranking cutoffs. This demonstrates its effectiveness in helping the recommendation model prioritize more relevant items, thereby better satisfying user preferences. The advantage of ConsRec can be attributed to its PLM-based ability to comprehend item text semantics and to capture item similarity patterns from user-item interaction signals. After training the recommendation module $M_{\text{Rec}}$ using denoised user-item interactions, ConsRec achieves substantial performance improvements, further highlighting its effectiveness.
5.4 The Effectiveness of Item Denoising in Enhancing the Recommendation Module
(a) TASTE.
(b) ConsRec.
Figure 5: Embedding Visualization of Items in Beauty. We randomly select a user and then use t-SNE to visualize item embeddings using TASTE and ConsRec.
@@ 5.4 The Effectiveness of Item Denoising in Enhancing the Recommendation Module
As illustrated in Figure 5, we investigate how the retained items contribute to improving the recommendation module during training. Specifically, we get the item representations retained and removed by $M_{\text{Filter}}$ using TASTE and ConsRec.
@@
Figure 5(a) shows the embedding space learned by TASTE, which is trained without item denoising. In this setting, the model yields a homogeneous embedding space, where retained and removed items are mixed together without clear separation. In contrast, Figure 5(b) demonstrates that ConsRec is able to better distinguish the removed items by mapping them to the periphery of the embedding space, effectively treating them as outliers. This observation highlights the effectiveness of ConsRec in leveraging denoising signals as prior knowledge to guide the recommendation module. By doing so, it facilitates a more discriminative representation space, where retained and removed items are clearly separated through contrastive training.
6 Conclusion
This paper introduces the User-Consistent Preference based Sequential Recommendation System (ConsRec), a novel framework designed to model user-consistent preferences for denoising user-interacted items. By explicitly filtering out noisy items, ConsRec enhances the interpretability of denoising and demonstrates its generalization ability, extending its advantages across various recommendation models.
Limitation
While ConsRec demonstrates effective denoising capabilities for sequential recommendations, the approach has several limitations that must be acknowledged. First, the performance of ConsRec is contingent on the quality of item descriptions; incomplete or low-quality descriptions may hinder the model’s ability to effectively filter out noisy user-item interactions. Secondly, for cold-start users with limited interaction history, ConsRec needs to build new user-item graphs from scratch, potentially introducing latency in generating personalized recommendations. This limitation could be mitigated by employing incremental graph updates, which dynamically add new edges as interactions occur, thereby reducing the need for full graph recomputation.
@@ Appendix A Appendix
A.1 License
We show the licenses of the datasets that we use. Yelp [^2] https://www.yelp.com/dataset uses Apache License 2.0, while Amazon Product shows its terms of use at website [^3] http://jmcauley.ucsd.edu/data/amazon/. All of these licenses and agreements allow their data for academic use.
A.2 Additional Experimental Details of Data Preprocessing
@@
Table 4: Statistics of the Datasets Used for Modeling the Sequential Recommendation Module ($M_{\text{Rec}}$).
@@
Table 5: Statistics of Training Data Used in the Filter Module ($M_{\text{Filter}}$). The data is used for optimizing item representations to estimate their similarities.
In our experiments, we employ Recbole Zhao et al. (2021) for preprocessing the data for sequential recommendation modeling, maintaining the same experimental settings as prior works Zhou et al. (2020); Xie et al. (2022); Liu et al. (2023); Chen et al. (2024). Specifically, each user-item interaction sequence $\{v_{1},\dots,v_{t}\}$ is processed using the leave-one-out strategy Xie et al. (2022); Chen et al. (2018); Sun et al. (2019); Zhou et al. (2020) to construct the training, development, and testing sets. In detail, the testing set predicts the $t$-th item $v_{t}$ based on the previously interacted items $v_{1:t-1}$ of the user, while the development set predicts the $t-1$-th item $v_{t-1}$ given the interactions $v_{1:t-2}$. The training set predicts $v_{i}$ using $v_{1:i-1}$, where $11<i<t-$ 1. Items and users with fewer than five interactions are filtered out, and all user-item interactions are treated as implicit feedback Chen et al. (2018); Sun et al. (2019); Zhou et al. (2020); Xie et al. (2022); Liu et al. (2023); Chen et al. (2024). The data statistics are shown in Table 4.
@@
Then, we construct the dataset for training the language model to represent items for estimating their similarities (Eq. 11). We start with the training and development sets used in sequential recommendation modeling and further process them for the Next Item Prediction (NIP) task. We first resample user-item interaction sequences from different datasets, namely Beauty, Sports, Toys, and Yelp. Then, we balance the number of data points to construct the training and development sets. This approach mitigates the risk of overfitting caused by data imbalance of different recommendation tasks. As shown in Table 5, the data statistics for both the training and development sets used in item representation learning are presented.
@@
Following previous studies Xie et al. (2022); Liu et al. (2023), we conduct evaluations in the full ranking testing scenario Dallmann et al. (2021); Krichene and Rendle (2020), where items are ranked across the entire item set rather than sampling subsets. This evaluation set mitigates the inconsistencies often observed in sampled evaluations, offering a more realistic and comprehensive assessment of recommendation performance.
A.3 Additional Details of Baseline Models
We reproduce all item-ID based models and DWSRec Zhang et al. (2024b) using implementations provided by RecBole Zhao et al. (2021), maintaining consistent settings across all models for fair comparison. For CoWPiRec and TASTE, we adhere to the default experimental settings as specified in their original papers. For HLLM, due to computational constraints, we ultilize TinyLlama-1.1B Zhang et al. (2024c) as the item-LLM component, and the user-LLM is implemented as a custom Llama model Touvron et al. (2023) with two hidden layers, trained from scratch using the following settings: learning rate of 1e-4, batch size of 2, training for 5 epochs, and a negative sampling ratio of 1024 negative samples per positive sample.
(a) SASRec
(b) BERT4Rec
(c) TASTE
Figure 6: Hyperparameter Sensitivity Analysis. We conduct experiments on the threshold hyperparameter, $\tau$, with results derived from the Amazon Beauty dataset using the SASRec 6(a), BERT4Rec 6(b), and TASTE 6(c) backbone models.
A.4 Hyperparameter Sensitivity Analysis
We investigate the sensitivity of the key hyperparameter $\tau$, which determines the similarity threshold during user-item interaction graph construction, as described in Sec. 3.2. We systematically vary $\tau$ from 0.2 to 0.8 while maintaining all other experimental settings. Figure 6 illustrates the performance across four evaluation metrics (Recall@10/20 and NDCG@10/20) for three backbone models: SASRec, Bert4Rec, and TASTE.
The results reveal a consistent pattern across all models and metrics. At lower $\tau$ values (0.2-0.4), the filtering mechanism is overly permissive, allowing to much noise in the interaction data and resulting in suboptimal performance. As $\tau$ increases to moderate values (0.5-0.6), we observe substantial performance improvements as irrelevant interactions are effectively filtered out. For instance, SASRec achieves peak R@10 performance at approximately $\tau=0.6$, with similar bell-shaped trends evident across other models and metrics. When $\tau$ exceeds 0.7, the filtering becomes excessively strict, potentially discarding valuable interaction signals and causing performance deterioration.
A.5 Ablation Study on Different Filter Models
To evaluate the effectiveness of the $M_{\text{Filter}}$ component in ConsRec, we conduct ablation studies comparing our approach with two alternative filter models–a vanilla T5 $M_{\text{Filter}}$ initialized from the T5-base checkpoint without additional training, and a TASTE $M_{\text{Filter}}$ trained specifically on recommendation data.
Dataset Method R@20 N@20
Beauty ConsRec 0.1774 0.0772
w/ T5 $M_{\text{Filter}}$ 0.1761 0.0762
w/ TASTE $M_{\text{Filter}}$ 0.1767 0.0767
Sports ConsRec 0.1082 0.0496
w/ T5 $M_{\text{Filter}}$ 0.1073 0.0457
w/ TASTE $M_{\text{Filter}}$ 0.1077 0.0471
Toys ConsRec 0.1896 0.0826
w/ T5 $M_{\text{Filter}}$ 0.1879 0.0802
w/ TASTE $M_{\text{Filter}}$ 0.1883 0.0811
Yelp ConsRec 0.1173 0.0509
w/ T5 $M_{\text{Filter}}$ 0.1169 0.0507
w/ TASTE $M_{\text{Filter}}$ 0.1170 0.0508
Table 6: Ablation Study on Different Filter Models in ConsRec.
As shown in Table 6, ConsRec consistently outperforms both alternative implementations across all datasets, demonstrating the effectiveness of our proposed training strategy for the filter module. The vanilla T5 $M_{\text{Filter}}$ exhibits the weakest performance, as it lacks domain-specific knowledge for recommendation tasks. The TASTE $M_{\text{Filter}}$ shows improved performance due to its training on recommendation tasks, but its effectiveness is limited by exposure to only a single dataset during training. In contrast, our ConsRec $M_{\text{Filter}}$ benefits from data-balanced pre-training across multiple datasets, resulting in more robust and generalizable item representations. This confirms that our training approach effectively enhances the model’s ability to capture relevant item characteristics while maintaining strong generalization capabilities across different recommendation scenarios.
Case #1 in Amazon Beauty
[…]
Description: … Valentina By Valentino 1.7 Oz Eau De Parfum Spray
[…]
Add Edge: $(n_{1},n_{2})$, $(n_{1},n_{4})$, $(n_{2},n_{4})$
The maximum connected subgraph $C_{max}$ is $\left\{1,2,4\right\}$, so that the filtered interaction sequence $\mathcal{H}_{\text{Filter}}$ is: $\left\{1,2,4\right\}$
Case #2 in Yelp
Node_ID: $n_{1}$    Address: Las Vegas, NV    Name: Maggiano’s Little Italy
Description:Nightlife, Restaurants, Italian, Bars …
[…]
Description:Nightlife, Food, Beer, Wine & Spirits, Pizza, Burgers …
[…]
Description:Veterinarians, Pets, Pet Services
[…]
Add Edge: $(n_{1},n_{2})$, $(n_{1},n_{4})$, $(n_{2},n_{4})$
The maximum connected subgraph $C_{max}$ is $\left\{1,2,4\right\}$, so that the filtered interaction sequence $\mathcal{H}_{\text{Filter}}$ is: $\left\{1,2,4\right\}$
Table 7: Case Studies. We present cases from two datasets, Amazon Beauty and Yelp. Different colors are highlighted to distinguish preference types: Blue represents majority preference, and Orange represents discrete interaction.
A.6 Case Studies
In Table 7, we present two examples from the Amazon Beauty and Yelp datasets to evaluate the effectiveness of ConsRec. For each dataset, a user is randomly selected, and we illustrate the process of denoising user-interacted items by modeling user-consistent preferences. In the case studies, we detail the procedure of identifying the maximum connected subgraph from the user-interacted item graph for item denoising.
As shown in the first case from the Amazon Beauty dataset, the user primarily interacts with face cream products focused on moisturizing, such as “Bare Escentuals bareMinerals Purely Nourishing Moisturizer” and “Ole Henriksen Truth Creme Advanced Hydration”. However, the user accidentally clicks on a product–“VALENTINO VALENTINA by Valentino for WOMEN: EAU DE PARFUM SPRAY 1.7 OZ”–due to their instant interests, which are significantly different in terms of both product types and attributes. In detail, while other items in the sequence emphasize “hydrating” and “moisturizing”, the “Parfum Spray” is unrelated to these preferences. When ConsRec constructs the user-interacted item graph, the “Parfum Spray” product fails to form edges with other items, indicating that this interaction is likely an outlier and should be excluded from user-interacted items. Consequently, the denoised item sequence $\mathcal{H}_{\text{Filter}}$ includes only products aligned with the consistent preferences of a specific user, such as moisturizing creams, enabling the recommendation module to better model user behavior from the user-item interactions.
In the second case, we present a specific user from the Yelp dataset who visits various locations. This user primarily interacts with places characterized by attributes such as “Nightlife”, “Restaurants”, and “Bars”, demonstrating a consistent pattern of preferences. However, one of the interacted locations– the “Spay & Neuter Center of Southern Nevada”–is a pet hospital, which is distinct from the user-consistent preference for nightlife and dining. Although this pet hospital is situated on the same street as other frequently visited locations, it likely represents an occasional visit rather than a reflection of the general behavior of the user. ConsRec assigns a lower similarity score between this noisy item and other user-interacted items, effectively excluding the noisy item from the sequence of user-interacted history. This enables the recommendation model to focus on more representative items, such as “Maggiano’s Little Italy” and “Weera Thai Restaurant”.
A.7 Theoretical Analysis of Consistent User Preference Modeling via ConsRec

### v2

@@ 5.4. Effectiveness of Different Item Representations for Shortcut Extraction
Figure 5. Effectiveness of LISRec in Learning Item Representations for Semantic Shortcut Extraction. Figure 5(a) and Figure 5(b) illustrate the item similarities among $M_{\text{Filter}}$-retained items and the ground truth items.
@@
As illustrated in Figure 5, we analyze the text similarity among the items retained by $M_{\text{Filter}}$ using the BGE-Base model to compute pairwise similarity scores. In Figure 5(a), we report the average similarity among the items selected by different semantic shortcut extraction strategies. The results show that all methods increase the internal similarity among selected items. Notably, when using LISRec, the extracted item subset achieves the highest average similarity, indicating stronger semantic alignment within the shortcut.
Furthermore, Figure 5(b) presents the average similarity between the $M_{\text{Filter}}$-retained user-interacted items and the ground-truth target item. Among all methods, LISRec yields the highest similarity, demonstrating its effectiveness in maintaining items that are semantically aligned with the target item that users need.
@@
As shown in Figure 4, we further evaluate the impact of different shortcut extraction methods on overall recommendation performance. Figures 4(a) and 4(b) report the performance of the recommendation module $M_{\text{Rec}}$ when trained on either the raw user-item interaction set $\mathcal{H}$ or the shortcut subset $\mathcal{H}_{\text{Shortcut}}$ extracted by $M_{\text{Filter}}$. In both cases, LISRec consistently outperforms other item representation models, with particularly strong gains at larger ranking cutoffs. This demonstrates its effectiveness in enabling the recommendation model to focus on the most relevant items, thereby better capturing stable user preferences. The advantage of LISRec is attributed to its PLM-based ability to capture item text semantics and to identify semantically meaningful patterns in user-item interactions. After training the recommendation module $M_{\text{Rec}}$ using the shortcut-extracted interaction histories, LISRec achieves substantial performance improvements, further highlighting the effectiveness of semantic shortcut extraction.
Table 5. Case Studies. We present cases from two datasets, Amazon Beauty and Yelp. Different colors are highlighted to distinguish preference types: Blue represents majority preference, and Orange represents discrete interaction.
Case #1 in Amazon Beauty
Node_ID: $n_{2}$      Name: Ole Henriksen Truth Creme Advanced Hydration, 1.7 Fluid Ounce
Node_ID: $n_{3}$      Name: VALENTINO VALENTINA by Valentino for WOMEN: EAU DE PARFUM SPRAY 1.7 OZ
Description: … Valentina By Valentino 1.7 Oz Eau De Parfum Spray
Node_ID: $n_{4}$      Name: Boscia Oil-Free Nightly Hydration, 1.4-Fluid Ounce
Add Edge: $(n_{1},n_{2})$, $(n_{1},n_{4})$, $(n_{2},n_{4})$
The maximum connected subgraph $C_{\text{max}}$ is $\left\{1,2,4\right\}$, so that the filtered interaction sequence $\mathcal{H}_{\text{Shortcut}}$ is: $\left\{1,2,4\right\}$
Case #2 in Yelp
Node_ID: $n_{1}$      Address: Las Vegas, NV      Name: Maggiano’s Little Italy
Description:Nightlife, Restaurants, Italian, Bars …
Node_ID: $n_{2}$      Address: Las Vegas, NV      Name: Lazy Dog Restaurant & Bar
Description:Nightlife, Food, Beer, Wine & Spirits, Pizza, Burgers …
Node_ID: $n_{3}$      Address: Las Vegas, NV      Name: Spay & Neuter Center of Southern Nevada
Description:Veterinarians, Pets, Pet Services
Node_ID: $n_{4}$      Address: Las Vegas, NV      Name: Weera Thai Restaurant - Sahara
Add Edge: $(n_{1},n_{2})$, $(n_{1},n_{4})$, $(n_{2},n_{4})$
The maximum connected subgraph $C_{\text{max}}$ is $\left\{1,2,4\right\}$, so that the filtered interaction sequence $\mathcal{H}_{\text{Shortcut}}$ is: $\left\{1,2,4\right\}$
(a) Similarity Scores between User Representation and Target Items.
(b) Embedding Visualization of Items using LISRec.
Figure 6. Effectiveness of Semantic Shortcut Extraction in Enhancing the Recommendation Module. Figure 6(a) and Figure 6(b) show the improvement of user and item representations after extracting shortcuts.
5.5. Effectiveness of Shortcut Extraction in Enhancing the Recommendation Module
@@
As illustrated in Figure 6, we investigate how the shortcut-retained items contribute to improving the recommendation module during training.
For each user, we first compute the embedding similarity between the user representation and the corresponding ground-truth item, under both the original and shortcut-applied training settings. Figure 6(a) shows that applying shortcut extraction during training consistently increases this similarity. This finding confirms that shortcut extraction strengthens the semantic alignment between user and target item representations, leading to more faithful modeling of user intent.
@@
Furthermore, we randomly select a user and then use t-SNE to visualize item embeddings. As shown in Figure 6(b), LISRec is able to better distinguish the removed items by mapping them to the periphery of the embedding space, effectively treating them as outliers. This observation highlights the effectiveness of LISRec in leveraging denoising signals as prior knowledge to guide the recommendation module. By doing so, it facilitates a more discriminative representation space, where retained and removed items are clearly separated through contrastive training.
5.6. Case Studies
In Table 5, we present two examples from the Amazon Beauty and Yelp datasets to evaluate the effectiveness of LISRec. For each dataset, we randomly select a user and showcase how semantic shortcut extraction operates on their interaction history. In the case studies, we detail the procedure of constructing the personalized item interaction graph and extracting the maximal preference-consistent subgraph, which serves to filter out noisy interactions.
As shown in the first case from the Amazon Beauty dataset, the user primarily interacts with face cream products focused on moisturizing, such as “Bare Escentuals bareMinerals Purely Nourishing Moisturizer” and “Ole Henriksen Truth Creme Advanced Hydration”. However, the user accidentally clicks on a product–“VALENTINO VALENTINA by Valentino for WOMEN: EAU DE PARFUM SPRAY 1.7 OZ”–due to their instant interests, which are significantly different in terms of both product types and attributes. In detail, while other items in the sequence emphasize “hydrating” and “moisturizing”, the “Parfum Spray” is unrelated to these preferences. When LISRec constructs the user-interacted item graph, the “Parfum Spray” product fails to form edges with other items, indicating that this interaction is likely an outlier and should be excluded from user-interacted items. Consequently, the denoised item sequence $\mathcal{H}_{\text{Shortcut}}$ includes only products aligned with the consistent preferences of a specific user, such as moisturizing creams, enabling the recommendation module to better model user behavior from the user-item interactions.
In the second case, we present a specific user from the Yelp dataset who visits various locations. This user primarily interacts with places characterized by attributes such as “Nightlife”, “Restaurants”, and “Bars”, demonstrating a consistent pattern of preferences. However, one of the interacted locations–the “Spay & Neuter Center of Southern Nevada”–is a pet hospital, which is distinct from the user-consistent preference for nightlife and dining. Although this pet hospital is situated on the same street as other frequently visited locations, it likely represents an occasional visit rather than a reflection of the general behavior of the user. LISRec assigns a lower similarity score between this noisy item and other user-interacted items, effectively excluding the noisy item from the sequence of user-interacted history. This enables the recommendation model to focus on more representative items, such as “Maggiano’s Little Italy” and “Weera Thai Restaurant”.
6. Conclusion
This paper introduces the Learned Item Shortcuts for Recommendation System (LISRec), a new framework that models stable user preferences by identifying dominant item subsets within a latent semantic space. LISRec leverages universal item representations to construct semantically meaningful shortcuts in user interaction histories. Our experiments on Yelp and Amazon Product datasets show that LISRec consistently outperforms baseline and previous text-based models. The results demonstrate that instruction-tuned universal item representations enable more robust shortcut selection and lead to more accurate, preference-consistent recommendations.
Acknowledgements.
This work is partly supported by the National Natural Science Foundation of China (No. 62206042 and No. 62461146205). This work is also supported by the AI9Stars community.
@@ Appendix A Appendix
A.1. License
We show the licenses of the datasets that we use. Yelp [^5] https://www.yelp.com/dataset uses Apache License 2.0, while Amazon Product shows its terms of use on website [^6] http://jmcauley.ucsd.edu/data/amazon/. All of these licenses and agreements allow their data for academic use.
A.2. Additional Experimental Details of Data Preprocessing
In our experiments, we employ Recbole (zhao2021recbole) for preprocessing the data for sequential recommendation modeling, maintaining the same experimental settings as prior works (Zhou2020s3; Xie2022DIF-SR; liu2023text; chen2024hllm). Specifically, each user-item interaction sequence $\{v_{1},\dots,v_{t}\}$ is processed using the leave-one-out strategy (Xie2022DIF-SR; chen2018sequential; Sun2019Bert4rec; Zhou2020s3) to construct the training, development, and testing sets. In detail, the testing set predicts the $t$-th item $v_{t}$ based on the previously interacted items $v_{1:t-1}$ of the user, while the development set predicts the t1t-1-th item $v_{t-1}$ given the interactions $v_{1:t-2}$. The training set predicts $v_{i}$ using $v_{1:i-1}$, where 1<i<t11<i<t-1. Items and users with fewer than five interactions are filtered out, and all user-item interactions are treated as implicit feedback (chen2018sequential; Sun2019Bert4rec; Zhou2020s3; Xie2022DIF-SR; liu2023text; chen2024hllm). The data statistics are shown in Table 6.
Table 6. Statistics of the Datasets Used for Modeling the Sequential Recommendation Module ($M_{\text{Rec}}$).
@@ A.2. Additional Experimental Details of Data Preprocessing
Table 7. Statistics of Training Data Used in the Filter Module ($M_{\text{Filter}}$). The data is used for optimizing item representations to estimate their similarities.
@@
Then, we construct the dataset for training the language model to represent items for estimating their similarities (Eq. 10). We start with the training and development sets used in sequential recommendation modeling and further process them for the Next Item Prediction (NIP) task. We first resample user-item interaction sequences from different datasets, namely Beauty, Sports, Toys, and Yelp. Then, we balance the number of data points to construct the training and development sets. This approach mitigates the risk of overfitting caused by data imbalance in different recommendation tasks.
As shown in Table 7, the data statistics for both the training and development sets used in item representation learning are presented.
@@
Following previous studies (Xie2022DIF-SR; liu2023text), we conduct evaluations in the full ranking testing scenario (dallmann2021case; krichene2022sampled), where items are ranked across the entire item set rather than sampling subsets. This evaluation set mitigates the inconsistencies often observed in sampled evaluations, offering a more realistic and comprehensive assessment of recommendation performance.
A.3. Theoretical Analysis of Stable User Preference Modeling via LISRec